Sketch the stream lines represented by ψ = xy. Also find out the velocity and its direction at point (2, 3).
1
Expert's answer
2020-12-10T17:26:26-0500

Velocity:

u=ui+vj


"u=\frac{\partial\psi}{\partial y}=x"


"v=-\frac{\partial\psi}{\partial x}=-y"


At point (2,3):

u=2

v=-3

u=2i-3j